OpenStudy (anonymous):

Solve on the interval {0,2Pi} (cosx+1)(2cos^2 x -3cosx -2) = 0 x=pi x=pi/6 x=pi/3 x=5pi/4

OpenStudy (amoodarya):

hint a= cos x (a+1)(2a^2-3a-2)=0

OpenStudy (amoodarya):

a=-1 a=(3+-root(9-4(2)(-2)))/4 so a=2, -1/2

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

@amoodarya